OpenStudy (jenniferjuice):

reward 5 medals What is the axis of symmetry for f(x) = 2x2 + 4x + 2? x = −3 x = −1 x = 1 x = 3

OpenStudy (anonymous):

lol...u can only give 1 medal at a time unless u have 5 accounts..:)

OpenStudy (jenniferjuice):

not true i use one account and ive given more than one medal

OpenStudy (anonymous):

lol...false

OpenStudy (jenniferjuice):

whoever answers this with the correct answer will tell you they got five medals when they get five medals

OpenStudy (jenniferjuice):

@bradely there is no -2 as an answer choice? is it -3?

OpenStudy (bradely):

f(x) = 2x^2 + 4x + 2 It is equation of parabola x=-b/2a a=2 ,b=4 x=-4/4 =-1 Axis of symmetry x=-1 option 2 is correct source: http://www.mathskey.com/question2answer/ask

OpenStudy (jenniferjuice):

@bradely i gave you a little over 5 but your welcome and thank you for helping c:

OpenStudy (bradely):

u r welcome